how many kg dinitrigen monoxide are present in 2.41 L of gas at STP? (answer as scientific notation)
1
Expert's answer
2020-12-21T03:38:41-0500

n=V1/V0= 2.4/22.4 = 0,107142857.

0,107142857 x 0.044 = 0,00471428571 kg